# Koby Hayashi

### v1

@@ 4.3.2 Analysis of Hybrid Sampling
which proves Lemma 4. ∎
The second Structural Condition for Hybrid Sampling is given by Lemma 9.
Teorema 5.
Let the thin SVD of ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times k}$ be ${\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}^{\sf T}$, let $0<\tau\leq 1$ be a threshold parameter with corresponding $\theta$, $0\leq\theta\leq k$, and define the nonnegative least squares problem ${\bm{\mathbf{{x}}}}_{nls}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}^{2}$ with ${\bm{\mathbf{{b}}}}\in\mathbb{R}^{m\times n}$. Let ${\bm{\mathbf{{S}}}}_{H}$ be the Hybrid Sampling matrix for ${\bm{\mathbf{{A}}}}$ with $\frac{2\xi}{\delta\epsilon_{r}}\leq s_{R}$. Then the following holds with at least probability $1-\delta$
@@
Proof.
Define ${\bm{\mathbf{{S}}}}_{\bar{D}}$ to be the $(m-s_{H}\times m)$ matrix that picks out the rows of ${\bm{\mathbf{{U}}}}_{A}$ that are not deterministically sampled by ${\bm{\mathbf{{S}}}}_{D}$. Row-partition the vector ${\bm{\mathbf{{r}}}}_{nls}$ conformally to ${\bm{\mathbf{{U}}}}_{A}$ as ${\bm{\mathbf{{r}}}}_{nls}=\begin{bmatrix}{\bm{\mathbf{{r}}}}_{D}\\ {\bm{\mathbf{{r}}}}_{R}\end{bmatrix}$. Then we have
@@
From this form we can apply Lemma 9 to obtain
@@
To succeed with probability at least $\delta$, we need $\frac{2\xi}{\epsilon_{r}\delta}\leq s_{R}$. ∎
Lemma 4 and Lemma 9 tell us that if $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$ proportion of the $k$ total leverage score ‘mass’ is sampled deterministically, one need only to take $\xi\phi$ samples, where $\xi=k-\theta$, instead of $k\phi$ to achieve the same theoretical guarantees for NLS or OLS problems, where $\phi=k\max(C\log(k/\delta),1/(\delta\epsilon_{r}))$ and $C=144/(1-\sqrt{2})^{2}$. Of course this relies on the deterministic samples accounting for a sufficiently large $\theta$. This helps to explain why the hybrid method typically outperforms the deterministic method when the same number of samples are taken and supports the experimental results for the CP decomposition of sparse tensors in [27].
@@ 5 Experimental Results
This section presents empirical findings on the proposed SymNMF algorithms. The section focuses on two data sets, one dense and one sparse. Both data sets are represented as graphs and the proposed SymNMF algorithms are used to cluster the vertices of the graphs as in previous works [19, 25]. For all experiments we use the initialization strategy from [25], however there are other initialization techniques such as those based on the SVD [2].
@@ 5.1 Web of Science Text Data
Once the algorithms are unable to reduce the normalized residual by more than 1e-4 for four consecutive iterations the methods stop. If Iterative Refinement (IR), Section 3.3, is being used the method will then switch over and use the same stopping criteria to determine when to stop iterating using the new method. All methods use Ada-RRF to determine how many power iterations to perform. Ada-RRF iterates until the normalized residual ceases to decrease by 1e-3 per power iteration.
The regularized version of SymNMF in Equation 3 requires an input hyperparameter $\alpha$ to balance the two objectives. Of course the compressed problem also needs this hyperparameter. In practice we find that using the value $\alpha=\max({\bm{\mathbf{{X}}}})$ recommended for the uncompressed problem works well [25].
@@ 5.1.1 Results
The convergence results for these experiments are shown in Figure 4. These show that the LAI-SymNMF method results in significant speed up. LAI-SymNMF without Iterative Refinement (IR) achieves about a 4$\times$ speed up over standard SymNMF with BPP. For LAI-SymNMF with HALS we observe a speed up of 7.5$\times$. For both BPP and HALS, IR does not appear to be needed when Ada-RRF finds a good starting low-rank approximation. For comparison we include the SymNMF variant of the “Compressed NMF” method, from [41], and see that it performs nearly identically to LAI-SymNMF. See Section A.2 for a discussion of the similarity of the two methods. In terms of the PGNCG method in Figure 3(c) and Table 2 we see that randomization greatly benefits PGNCG resulting in about a 6×6$\times$6 × speed up. Additionally, IR in conjunction with PGNCG achieves the lowest normalized residual norm. Overall we see that randomized methods for SymNMF result in significant speed up while preserving, or even improving, solution quality. Note that we do not report results for LvS-SymNMF on this data set as the code took prohibitively long to execute.
Plots showing the projected gradient values for each algorithm can be found in Section C.2. We include this metric as it is often used as an alternative stopping criteria for NMF and its variants [23, 13]. Plots and tables showing results for different values of the oversampling parameter $\rho$ are also provided. We find that varying $\rho$ does not have much effect on the solution quality but can negatively impact run time. Finally, Section C.2 includes plots and tables showing that the use of Algorithm Ada-RRF provides better overall results than using a static choice of $q=2$ with IR.
We briefly compare against Spectral Clustering as described in [35, 19] to validate SymNMF-based clustering results. Spectral clustering achieves an average ARI of $0.293$ over 20 runs. This is a worse ARI score than all of the SymNMF methods we try, see Table 2. We also note that the normalized residual achieved by a rank $=7$ SVD is $0.9340$.
@@
Table 2: Metrics for the various run shown in fig. 4. Iters is mean number of iterations taken, Tim is mean run time in seconds, Avg. Min-Res is average minimum achieved residual, Min-res is overall lowest achieved residual, and Mean-ARI is the mean ARI score. Averages are taken over 20 runs.
@@ 5.2 Microsoft Open Academic Graph
The regularized version of SymNMF in Equation 3 requires an input hyperparamter $\alpha$ to balance the two objectives. Of course the sampled problem will also need this hyperparameter. We propose that using the same values of $\alpha$ for the leverage score sampling method and the deterministic method is reasonable since $\mathbb{E}[\|{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{H}}}}\|_{F}^{2}]=\|{\bm{\mathbf{{H}}}}\|_{F}^{2}$ and $\mathbb{E}[\|{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}}\|_{F}^{2}]=\|{\bm{\mathbf{{W}}}}\|_{F}^{2}$.
@@
Algorithm LvS-SymNMF is run with BPP and HALS as the update rules. The normalized residual norms and the projected gradient values are shown in Figure 2. First we note that the LAI method with BPP update rule is generally unsuccessful at reducing the error. We hypothesize that this is because the RRF produces a LAI with a large value of $\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}-{\bm{\mathbf{{W}}}}_{+}({\bm{\mathbf{{H}}}}_{+})^{\sf T}\|_{F}$ as in Equation 23.
@@
Next we note that using the Hybrid Sampling method is important to achieve good speed-up. In the case of the BPP update rule, in Figure 1(a) one can observe that the purely random method ($\tau=1$) does not provide any speed up over the standard BPP method. However when the hybrid method with $\tau=1/s$ is used one can see that the leverage score sampling algorithm becomes competitive. When the BPP update rule is used, a significant amount of the time per iteration goes into solving the $2m$ NLS problems. So while leverage score sampling effectively eliminates the cost of multiplying by ${\bm{\mathbf{{X}}}}$, only about a $50\%$ speed up is achieved as leverage score sampling has no effect on the cost of the update function. In the case of HALS, we observe significantly better speed up when leverage score sampling is used. Figure 1(a) clearly shows that hybrid leverage score sampling outperforms both leverage score sampling and the standard HALS method. The hybrid leverage score sampling method in this case is able to achieve about a $\times5$ speed up per iteration over standard HALS. We note that the use of our new HALS update rules, as in Equations 13 and 14, is crucial for obtaining good performance.
@@
Figure 3 shows the time break down for 3 algorithms: leverage score sampling with HALS and BPP as the update rule and a standard NMF method with HALS as the update rule. The timings are divided into 3 categories: 1) “Matrix Multiplication” (MM) for computing the four main matrix products, 2) “Solve” for the time spent in the update rule (e.g. time spent in the BPP solver), and 3) “Sampling” time spent performing leverage score sampling which includes computing the thin QR decomposition. We can see that the MM time is greatly reduced by using leverage score sampling while incurring an acceptable overhead in terms of sampling time. Lastly we see that this data confirms that the BPP solver is limiting potential computational gains from leverage score sampling.
@@
Figure 2: Normalized residual and projected gradient values for OAG using HALS and BPP update rules. $k=16$, $\tau=1/s$.
@@ 5.2.1 Results for the OAG
We now analyze the output of running the LvS-SymNMF algorithms on the OAG. The factors with $k=16$ are used to form $k$ clusters, as in the experiment before on the WoS hypergraph, following the methodology in [25]. Since this data set does not come with ground truth, assessing the quality of the LvS-SymNMF outputs is more difficult. The residuals output by both the HALS and the LvS-HALS method differ by a tiny amount, on the order of 1e-12, with the leverage score method achieving the lower residual.
@@
Silhouette Scores range from $1$ to $-1$, with $1$ meaning a vertex is very similar to its own cluster and $-1$ meaning it should be moved to its next closest cluster. Individual vertex scores are then averaged over their cluster memberships to obtain cluster level Silhouette Scores. Note that this definition of Silhouette Score is for similarity metrics and so is slightly different that the typical Silhouette Score equation which is defined for dissimilarity metrics.
The Silhouette Scores are computed for each cluster. The Silhouette Scores for the HALS methods are perfect with the exception of Cluster-0, which still achieves a high score of 0.78. The Silhouette Scores for the LVS-HALS algorithm are not as high as the ones for HALS but are still well above 0. The exception is Cluster-0 (from LVS-HALS), which has a near 0 Silhouette Score of $-0.04$, meaning that it does not form a good cluster but also that its vertices do not necessarily belong in one of the other clusters. Other than Cluster-0 the LVS-HALS method finds clusters with mean Silhouette Score $0.92\pm 0.1$ ($\pm$ Standard Deviation). This indicates that the LVS-HALS method is able to find well structured clusters in terms of Silhouette Score. To further validate that LVS-HALS is finding meaningful clusters we show the top 10 words associated with Clusters 10 to 15 in Section 5.2.1. More data about this experiment can be found in Section D.2.

### v2

@@ 4.3.2 Analysis of Hybrid Sampling
which proves Lemma 4.3.
The second Structural Condition for Hybrid Sampling is given by Lemma 4.5.
Lemma 4.5.
Let the thin SVD of ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times k}$ be ${\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}^{\sf T}$ and define the nonnegative least squares problem ${\bm{\mathbf{{x}}}}_{nls}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}^{2}$ with ${\bm{\mathbf{{b}}}}\in\mathbb{R}^{m\times n}$. Let $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$ be the sum of the leverage scores corresponding to rows in the deterministically included set. Let ${\bm{\mathbf{{S}}}}_{H}$ be the Hybrid Sampling matrix for ${\bm{\mathbf{{A}}}}$ with $\frac{2\xi}{\delta\epsilon_{r}}\leq s_{R}$. Then the following holds with probability at least $1-\delta$
@@
Proof 4.6.
Define ${\bm{\mathbf{{S}}}}_{\bar{D}}$ to be the $(m-s_{H})\times m$ matrix that picks out the rows of ${\bm{\mathbf{{U}}}}_{A}$ that are not deterministically sampled by ${\bm{\mathbf{{S}}}}_{D}$. Row-partition the vector ${\bm{\mathbf{{r}}}}_{nls}$ conformally to ${\bm{\mathbf{{U}}}}_{A}$ as ${\bm{\mathbf{{r}}}}_{nls}=\begin{bmatrix}{\bm{\mathbf{{r}}}}_{D}\\ {\bm{\mathbf{{r}}}}_{R}\end{bmatrix}$. Then we have
@@
From this form we can apply Lemma F.3 to obtain
@@
To succeed with probability at least $\delta$, we need $\frac{2\xi}{\epsilon_{r}\delta}\leq s_{R}$.
Discussion of Lemma 4.3 and Lemma 4.5
Lemma 4.3 and Lemma 4.5 tell us that if $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$ proportion of the $k$ total leverage score ‘mass’ is taken deterministically with $s_{D}$ rows, one needs only to take an additional $\xi\phi$ random samples, where $\xi=k-\theta$, instead of $k\phi$ to achieve the same theoretical guarantees for NLS or OLS problems, where $\phi=\max(C\log(k/\delta),1/(\delta\epsilon_{r}))$ and $C=144/(1-\sqrt{2})^{2}$. Overall hybrid sampling requires $s_{D}+\xi\phi$ samples whereas standard leverage score sampling requires $k\phi$ samples. For hybrid sampling to result in a lower sample complexity we must have $s_{D}<(k-\xi)\phi=\theta\phi$. This relies on the deterministic samples accounting for a sufficiently large $\theta$. This helps to explain why the hybrid method typically outperforms the deterministic method when the same number of samples are taken and supports the experimental results for the CP decomposition of sparse tensors in [36].
To determine the deterministic inclusion set $\mathscr{I}_{D}$ we use the thresholding technique from [36]. This method chooses a parameter $0<\tau\leq 1$ and then the deterministic set includes all rows such that $\tau\leq l_{i}({\bm{\mathbf{{A}}}})/k$. This means that all rows with a leverage score value greater or equal to the threshold $\tau$ are deterministically included. Lemma 4.3 and Lemma 4.5 do not strictly rely on using a thresholding technique to determine which rows should be deterministically included. Both results easily hold in the case that a different method is used to determine the set $\mathscr{I}_{D}$. Any such technique for determining $\mathscr{I}_{D}$, like the thresholding one, will likely want to keep $|s_{D}|$ small while making $\theta$ as large as possible.
@@ 5 Experimental Results
This section presents empirical findings on the proposed SymNMF algorithms. The section focuses on two data sets, one dense and one sparse. Both data sets are represented as graphs and the proposed SymNMF algorithms are used to cluster the vertices of the graphs following approaches given in previous work [27, 35]. Namely, clustering via [27] proceeds by constructing a hypergraph random walk from the data, which is then used to construct a symmetric adjacency matrix that serves as the input to SymNMF clustering. For all experiments we use the initialization strategy from [35], though we acknowledge other initialization techniques such as those based on the SVD [5]. The strategy in [35] proceeds by generating a initial matrix for ${\bm{\mathbf{{H}}}}\in\mathbb{R}_{+}^{m\times k}$ with elements drawn uniformly from the interval on 0 to 1, and scaling them by the value $2*\sqrt{\zeta/k}$ where $\zeta$ is the average of all elements of ${\bm{\mathbf{{X}}}}$. Intuitively this ensures that the norm of the initial factor matrices is not too large or small in comparison to ${\bm{\mathbf{{X}}}}$. Given an output factor ${\bm{\mathbf{{H}}}}$ the cluster membership for the $i$th vertex $v$ is determined by finding the column index with the maximum value in the $i$th row of ${\bm{\mathbf{{H}}}}$ [35].
For checking convergence we use the normalized residual and the norm of the projected gradient. These are two standard metrics for checking the convergence of NMF and its variants [32, 19]. Definitions and some more discussion of these two metrics can be found in Appendix C.
@@ 5.1 Web of Science Text Data
Once the algorithms are unable to reduce the normalized residual by more than 1e-4 for four consecutive iterations the methods stop. If Iterative Refinement (IR), described in Section 3.3, is being used the method will switch over to using the full input data matrix and apply the same stopping criteria to determine when to stop fitting to the full input matrix. All methods use Ada-RRF to determine how many power iterations to perform. Ada-RRF iterates until the normalized residual ceases to decrease by 1e-3 per power iteration. Each method is run 10 times.
We use a labeling system to name the algorithms. The algorithms we consider here vary in terms of Update rules among 1) HALS, 2) BPP, 3) PGNCG, if they use LAI, Compression (Comp) as in [51], or neither, and finally if Iterative Refinement is used, denoted with IR. A combination of these labels indicates the method used. For example LAI-BPP-IR is the Algorithm LAI-SymNMF with BPP as the update rule and iterative refinement run at the end. LAI-BPP indicates the same technique but without IR at the end. Finally BPP indicates standard SymNMF method with BPP as the update rule.
The regularized version of SymNMF in Equation 3 requires an input hyperparameter $\alpha$ to balance the two objectives. The compressed problem also needs this hyperparameter. In practice we find that using the value $\alpha=\max({\bm{\mathbf{{X}}}})$ recommended for the uncompressed problem works well [35].
@@ 5.1.1 Results
The convergence results for these experiments are shown in Fig. 1. These show that the LAI-SymNMF method results in significant speed up. LAI-SymNMF without Iterative Refinement (IR) achieves about a 4$\times$ speed up over standard SymNMF with BPP. For LAI-SymNMF with HALS we observe a speed up of 7.5$\times$. For both BPP and HALS, IR does not appear to be needed when Ada-RRF finds a good starting low-rank approximation. For comparison we include the SymNMF variant of the “Compressed NMF” method, from [51], and see that it performs nearly identically to LAI-SymNMF. See Section B.1 for a discussion of the similarity of the two methods. In terms of the PGNCG method in Fig. 4(c) and Table 2 we see that randomization greatly benefits PGNCG resulting in about a 6×6$\times$6 × speed up. Additionally, IR in conjunction with PGNCG achieves the lowest normalized residual norm. Overall we see that randomized methods for SymNMF result in significant speed up while preserving, or even improving, solution quality. The run time of the randomized methods included the time needed to compute the LAI. This is why all the randomized methods appear to ‘start’ later than the non-randomized methods.
Note that we do not report results for LvS-SymNMF on this data set as the code took too long to execute. This is due to the fact that copying large portions (sampled rows) of a large dense data matrix ${\bm{\mathbf{{X}}}}$ at each iteration takes a long time.
Additional experimental data can be found in Section G.1. Fig. 4 and Tables 4 and 5 show results for different values of the oversampling parameter $\rho$. We find that varying $\rho$ does not have much effect on the solution quality but can negatively impact run time. Finally, Section G.1 includes plots and tables showing statistics for using a static choice of $q=2$ with IR. Comparing these results with those from using Algorithm Ada-RRF shows the superiority of Algorithm Ada-RRF versus a static choice of $q$.
We briefly compare against Spectral Clustering as described in [45, 27] to validate SymNMF-based clustering results. Spectral clustering achieves an average ARI of $0.293$ over 20 runs. This is a worse ARI score than all of the SymNMF methods we try, see Table 2. Spectral clustering using the MATLAB functions eigs() and kmeans() takes about $35$ seconds to execute. We also note that the normalized residual achieved by a rank $=7$ SVD is $0.9340$.
@@
Table 2: Metrics for the various run shown in Fig. 1. Iters is mean number of iterations taken, Time is mean run time in seconds, Avg. Min-Res is average minimum achieved residual, Min-res is overall lowest achieved residual, and Mean-ARI is the mean ARI score. Averages are taken over 20 runs.
@@ 5.2 Microsoft Open Academic Graph
The regularized version of SymNMF in Equation 3 requires an input hyperparameter $\alpha$ to balance the two objectives. The sampled problem will also need this hyperparameter. We propose that using the same values of $\alpha$ for the leverage score sampling method and the deterministic method is reasonable since $\mathbb{E}[\|{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{H}}}}\|_{F}^{2}]=\|{\bm{\mathbf{{H}}}}\|_{F}^{2}$ and $\mathbb{E}[\|{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{W}}}}\|_{F}^{2}]=\|{\bm{\mathbf{{W}}}}\|_{F}^{2}$.
@@
Algorithm LvS-SymNMF is run with BPP and HALS as the update rules. The normalized residual norms and the projected gradient values are shown in Fig. 2. See Section C.3 for details on projected gradient values. First we note that the LAI method with BPP update rule is generally unsuccessful at reducing the error. We hypothesize that this is because the RRF produces a LAI with a large value of $\|{\bm{\mathbf{{U}}}}_{X}{\bm{\mathbf{{V}}}}_{X}-{\bm{\mathbf{{W}}}}_{+}({\bm{\mathbf{{H}}}}_{+})^{\sf T}\|_{F}$ as in Equation 17.
@@
Next we note that using the Hybrid Sampling method is important to achieve good speed-up. In the case of the BPP update rule, in Fig. 2(a) one can observe that the purely random method ($\tau=1$) does not provide any speed up over the standard BPP method. However when the hybrid method with $\tau=1/s$ is used, one can see that the leverage score sampling algorithm becomes competitive. When the BPP update rule is used, a significant amount of the time per iteration goes into solving the $2m$ NLS problems. So while leverage score sampling effectively eliminates the cost of multiplying by ${\bm{\mathbf{{X}}}}$, only about a $50\%$ speed up is achieved as leverage score sampling has no effect on the cost of the update function. In the case of HALS, we observe significantly better speed up when leverage score sampling is used. Fig. 2(a) clearly shows that hybrid leverage score sampling outperforms both leverage score sampling and the standard HALS method. The hybrid leverage score sampling method in this case is able to achieve about a $\times5$ speed up per iteration over standard HALS. We note that the use of the modified HALS update rules, as in Equations 5 and 6, are crucial for obtaining good performance.
@@
Fig. 3 shows the time break down for 3 algorithms: leverage score sampling with HALS and BPP as the update rule and a standard NMF method with HALS as the update rule. The timings are divided into 3 categories: 1) “Matrix Multiplication” (MM) for computing the four main matrix products, 2) “Solve” for the time spent in the update rule (e.g. time spent in the BPP solver), and 3) “Sampling” time spent performing leverage score sampling which includes computing the thin QR decomposition. The sampling time also includes time to form the needed subsampled matrices. The OAG data set is stored using MATLAB’s sparse matrix representation which allows for fast row/column slicing (because the matrix is symmetric) via the Compressed Sparse Column or (CSC) format. This is in juxtaposition to the hypergraph representation of the WoS data set which is dense and so does not benefit from sparse matrix formats. Additionally, the MM time is greatly reduced by using leverage score sampling while incurring an acceptable overhead in terms of sampling time. Lastly we see that this data confirms that the BPP solver has limited potential computational gains from leverage score sampling.
Figure 2 also shows the projected gradient values for the OAG runs. As discussed in the next section, the solutions found by the randomized and non-randomized methods are quite different. The fact that the randomized methods does a good job of reducing both the Residual and the Projected Gradient helps increase confidence that Algorithm LvS-SymNMF is performing well. See Section C.3 for details on the Projected Gradient.
@@
Figure 2: Normalized residual and projected gradient values for OAG using HALS and BPP update rules. $k=16$, $\tau=1/s$ or $\tau=1$ as indicated in the legend.
@@ 5.2.1 Results for the OAG
We now analyze the output of running the LvS-SymNMF algorithms on the OAG. The factors with $k=16$ are used to form $k$ clusters, as in the WoS hypergraph, following the methodology in [35]. Since this data set does not come with ground truth, assessing the quality of the LvS-SymNMF outputs is more difficult. The residuals output by both the HALS and the LvS-HALS method differ by a tiny amount, on the order of 1e-12, with the leverage score method achieving the lower residual.
We use a similar labeling scheme as for the WoS experiments. LvS represents the use of leverage scores sampling and the considered update rules are HALS and BPP. So LvS-BPP represents Algorithm LvS-SymNMF with BPP. Additionally for leverage score sampling methods the value of $\tau$ used is also indicated.
@@
Silhouette Scores range from $1$ to $-1$, with $1$ meaning a vertex is very similar to its own cluster and $-1$ meaning it should be moved to its next closest cluster. Individual vertex scores are then averaged over their cluster memberships to obtain cluster level Silhouette Scores. Note that this definition of Silhouette Score is for similarity metrics and so is different that the typical Silhouette Score equation which is defined for dissimilarity metrics.
The Silhouette Scores are computed for each cluster. The Silhouette Scores for the HALS methods are perfect with the exception of Cluster-0, which still achieves a high score of 0.78. The Silhouette Scores for the LVS-HALS algorithm are not as high as the ones for HALS but are still well above 0. The exception is Cluster-0 (from LVS-HALS), which has a near 0 Silhouette Score of $-0.04$, meaning that it does not form a good cluster but also that its vertices do not necessarily belong in one of the other clusters. Other than Cluster-0 the LVS-HALS method finds clusters with mean Silhouette Score $0.92\pm 0.1$ ($\pm$ Standard Deviation). This indicates that the LVS-HALS method is able to find well structured clusters in terms of Silhouette Score. To further validate that LVS-HALS is finding meaningful clusters we show the top 10 words associated with Clusters 10 to 15 in Table 3. More data about this experiment can be found in Section G.2.
[…]